# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

### v1

@@ 4.2 National Production of Career Elites
The above analysis examines the distributional consequence of the centralized admissions. A complementary problem is its overall impact for the whole country. To study it, we now turn to whether the meritocratic centralization improved the national production of career elites in the long run. We focus on a specific group of elites — higher civil officials — for whom we have complete-count data from administrative records. Top-ranking civil officials were also the most relevant group of career elites; in our empirical context, higher civil service was considered to be one of the most prestigious jobs due to their political and legislative influence. As a result, high-achieving students at the top universities competed to get jobs at the Ministry of Finance and other selective ministries (Shimizu, 2019). We investigate whether cohorts exposed to Capp produced a greater national number of top-ranking civil officials compared to cohorts exposed to Dapp.
@@
Our main data source is the list of individuals who passed the Higher Civil Service Examinations (HCSE) and their biographical information (Hata, 1981). The HCSE were selective national qualification exams held annually from 1894 to 1947. The 1893 ordinance required all individuals to pass the HCSE for appointment in the administrative division of higher civil service with some exceptions for special appointments (Spaulding, 1967, Chapter 25; Shimizu, 2019, Chapter 5). We digitized the information of all individuals who passed the administrative division of the HCSE in 1894–1941, including their full name, education, year of university graduation, year of passing the exam, starting position, final position, year of retirement, and other notable positions held. Because education includes not only final but also the second to final education, we observe both university and higher school (if applicable).
In the following analysis, we define “top-ranking officials” as higher civil officials who were internally promoted to one of the top three ranks by the end of their career. In the prewar Japanese bureaucracy system, the higher civil service refers to the top ten ranks of national government offices in the administrative, judicial, and diplomatic divisions. Within the higher civil service, the top three ranks were distinctively called “imperial appointees.” The first rank consisted of minister level positions, while the second and third ranks consisted of vice minister level positions such as vice minister, director general, bureau chief, and prefectural governor. Appendix A.4.2 provides more details.
We count the number of top-ranking officials and the number of individuals who passed the administrative division of the HCSE (hereafter “exam passers”) produced by cohort. Here cohort is defined by “the year of entering a higher school or its equivalent.” Since we only observe the year of university graduation in the above data, we estimate “the year of entering a higher school or its equivalent” using the method described in Appendix A.4.2. We first find the exact year of entering a higher school for all top-ranking officials who graduated from Schools 1–8 by searching their names in Student Registers of Schools 1–8. Using this information, we then estimate “the year of entering a higher school or its equivalent” for the rest of higher civil officials.
@@
Effect on Higher Civil Service Exams Passers
As an intermediate outcome, we first examine the impacts of the centralized admission system on the number of individuals who passed the selective HCSE. We divide exam passers into three mutually exclusive subgroups: (a) those who graduated from School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates. For each subgroup, we count the number of exam passers by cohort. We estimate the following equation for the entire group and for each subgroup:
@@
where $Y_{t}$ is the number of exam passers in a given group in cohort $t$ (defined by the year of entering a higher school or its equivalent), and $Centralized_{t}$ is the indicator that takes 1 if cohort $t$ entered a higher school or its equivalent during Capp. For a subgroup regression, we control for the total number of exam passers in cohort $t$ (denoted by $X_{t}$). We also control for a quadratic time trend where $Trend_{t}$ is the number of years since 1897.
Capp changed the composition of exam passers. In Table 4 Panel (a), Column (3) shows that Capp increased the number of exam passers who graduated from Schools 2–8 by 23%, while Column (4) indicates that Capp reduced the number of exam passers who are not Schools 1–8 graduates by 13%. By contrast, Capp did not have significant effect on the total number of exam passers or the number of exam passers who graduated from School 1 (Columns (1) and (2)). In other words, the centralization had little impact on the total number of exam passers, but had a major impact on the composition of exam passers, and its positive effect was concentrated on Schools 2–8 graduates. This finding is consistent with the result of our short-run analysis that the centralization led to a large increase in Tokyo-area-born high-achieving entrants to Schools 2–8 (Table 1 Columns (3)–(9)), which likely improved the academic standing of these schools.
@@
Finally, we analyze the impact of the centralization on the national production of top-ranking officials. Our main outcome variable is the number of higher civil officials who were internally promoted to the top three ranks by the end of their career. We count the number of top-ranking officials by cohort and run the same regression as above for the entire group and for each of the following subgroups: (a) those who graduated from School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates. For all regressions, we control for the total number of exam passers from the cohort.
@@
Meritocratic centralization produced a larger number of high-quality bureaucratic elites for the whole country. Table 4 Panel (b) Column (1) shows that the total number of top-ranking officials increased for Capp cohorts by 15%, with high statistical significance. This result is consistent with the idea that the centralized system (which assigned top-scoring students to Schools 1-8) increased the quality of Schools 1–8 entrants, which in turn improved the likelihood of Schools 1–8 graduates to win promotion competitions over close Dapp cohorts. Indeed, Column (3) indicates that Capp had a large, positive, and significant effect on the number of top-ranking officials who graduated from Schools 2–8, where the number of urban-born high-achieving entrants spiked during centralization. The number of top-ranking officials who graduated from Schools 2–8 increased by 47% for Capp cohorts compared to Dapp cohorts. By contrast, in Column (2), Capp only had a small and insignificant effect on the number of top-ranking officials who graduated from School 1, where the number of urban-born high-scoring entrants did not change much during centralization.
@@
Schools 1-8’s impacts on the likelihood of promotion to top-ranking officials appear especially large for high-achieving students. Indeed, Capp’s positive effect on Schools 2–8 graduates was so large (Column (3)) that it dominated Capp’s small and negative effect on those who did not graduate from Schools 1–8 (Column (4)), resulting in a positive effect on the total number of top-ranking officials (Column (1)). This finding suggests that the meritocratic assignment of high achievers to Schools 1-8 produced a net positive effect on the national production of top-ranking officials. This result is inconsistent with the pure selection hypothesis which argues that the role of the centralized admissions is simply to select and send a fixed number of high-achieving students to national higher education and that national higher education itself does not give students any value added.
@@
Schools 1-8’s particularly large value-added for high-achieving students can be due to multiple factors. First, human capital returns to education by these schools could be higher for high-achieving students than lower-achieving students (due to demanding curriculum, for example). Gathering high-achieving students in these schools may also produce higher peer effects. Second, high-achieving students may benefit more from their signaling value or from gaining connections with powerful alumni than low-achieving students. While our data and variation do not allow for a clear decomposition into these different factors, our analysis suggests that the effect of meritocratic centralization remains similar even after controlling for the size of same-cohort colleagues from the same schools or focusing on those with the highest educational qualification (Appendix Section A.4.4). Much of Schools 1-8’s effects therefore likely come from the first human capital factor.
@@
One potential threat to our identification is that the number of available top-ranking positions might increase during the Capp periods. To investigate this concern, recall that we define our cohort by the year of entering a higher school or its equivalent (not by the year of becoming top-ranking officials). Within each cohort, the number of years taken from entering a higher school to the appointment for the first top-ranking position varied widely from 20 to 30 years (see Appendix Figure A.9 and Appendix A.4.3 for details). Individuals in a given cohort were therefore promoted to top-ranking positions in different years. This limits the concern that a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions may drive our results.
@@
Another potential threat is that the number of top-ranking officials may be differentially impacted by wars between Dapp and Capp cohorts. In particular, Japan was occupied by the US after WWII. During the postwar US Occupation, a substantial number of top officials were purged from public service, and a few were sentenced to death for war crimes. However, we show that the number of top-ranking officials who died in wars, or were purged or executed after WWII, changes little between Dapp and Capp cohorts (Appendix Table A.16).
@@
For career bureaucrats, the meritocratic centralization turns out to be more productively efficient than the decentralized system, as the government envisioned. Whether or not this was true for all career elites is an open question. For example, it is possible that the quality of career bureaucrats improved at the expense of other types of elites. To examine this possibility, we use the 1939 PIR data to compute the share of top bureaucrats among all “socially distinguished individuals” listed in the PIR. [^30] Top bureaucrats are defined as high-ranking central government officials listed in the PIR as in Section As Appendix Figure A.10 shows, we observe no positive association between the share of government officials and the cohort’s exposure to the centralized admissions (although the share of government officials is rising over time reflecting a growing public sector). Our result is unlikely to be driven by students’ brain drain from other elite occupations into top bureaucrats under the meritocratic centralization.
Finally, we obtain suggestive evidence that the meritocratic centralization might produce a greater number of elites in general. As Appendix Figure A.11 and Table A.17 show, higher exposure to the centralized system is associated with a greater total number of individuals listed in the PIR. A potential concern with this analysis is that this may be driven by the age effect. For example, people of certain ages may have had a higher chance of holding important positions and being selected into the PIR, while such ages may happen to be correlated with exposure to centralization. To distinguish the cohort effect from the age effect, we use both 1934 and 1939 editions of the PIR while controlling for the age of each cohort. The resulting estimates show that the number of individuals listed in the PIR is 4–5% higher among Capp cohorts. The centralized system may therefore be more productively efficient not only for producing top bureaucrats but also for producing other types of career elites.
@@ 5 Conclusion
The design of school admissions persistently impacts the production and geography of career elites. We reveal this fact by looking at the world’s first recorded use of nationally centralized admissions and its subsequent abolitions in early twentieth-century. While centralization was designed to make the school seat allocations more meritocratic, there turns out to be a tradeoff between meritocracy and equal regional access to elite higher education and career advancement. The meritocratic centralization led students to apply to more selective schools and make more inter-regional applications. As high-achieving students were located disproportionately in urban areas, however, centralization caused urban applicants to crowd out rural applicants from advancing to higher education.
@@
Most importantly, these impacts persist: Several decades later, the centralized system produced more top-ranking government officials, indicating that the meritocratic centralization may be more productively efficient. The distributional effect also persists: The meritocratic centralization increased the number of high income earners, medal recipients, and other elite professionals born in urban areas relative to those born in rural areas.
@@
Even though our study uses the admission reforms unique to our historical setting, the implications of our study may have broader relevance. For instance, distributional consequences of centralized meritocratic admissions may be a reason why many countries continue to use seemingly inefficient decentralized college admissions. While this paper focuses on the decentralized system with a single application, the takeaway from our results would be relevant for other types of decentralized admissions as long as costly applications make applicants self-select into a limited number of schools. For any such decentralized systems, the meritocracy-equity tradeoff would be an important potential concern.
@@
Methodologically, the use of natural experiments in history may be also valuable for studying the long-run effects of market designs in other areas, such as housing, labor, and health markets. It is the multiple bidirectional policy changes in history that allow us to measure the long-run effects. The disadvantage of using historical events is the limited availability of data. The ideal way to alleviate the data concerns would be to use modern administrative data. For example, one may imagine linking administrative tax return data and school district data to measure the long-run effects of school choice reforms around the world in the past few decades. Such an effort would be a fruitful complement to our historical study.
A. Baseline Specification
Age 17 under centralization 3.18 0.61 1.68 2.82 1.84 0.84 0.41
  $\times$ Tokyo area ($<$100 km) (0.027)** (0.016)** (0.060)* (0.010)*** (0.088)* (0.002)*** (0.025)**
[0.000]*** [0.020]** [0.008]*** [0.001]*** [0.045]** [0.016]** [0.182]
B. With Control Variables
Age 17 under centralization 1.93 0.52 1.59 2.44 1.57 0.63 0.39
  $\times$ Tokyo area ($<$100 km) (0.002)*** (0.025)** (0.026)** (0.002)*** (0.071)* (0.001)*** (0.041)**
[0.014]** [0.017]** [0.007]*** [0.000]*** [0.040]** [0.041]** [0.143]
C. Bidirectional Specification with Control Variables
[…]
  $\times$ Tokyo area ($<$100 km) (0.165) (0.078)* (0.006)*** (0.004)*** (0.057)* (0.010)*** (0.149)
[0.034]** [0.049]** [0.006]*** [0.003]*** [0.023]** [0.170] [0.359]
[…]
  $\times$ Tokyo area ($<$100 km) (0.001)*** (0.022)** (0.183) (0.004)*** (0.338) (0.006)*** (0.025)**
[0.005]*** [0.131] [0.021]** [0.008]*** [0.488] [0.032]** [0.079]*
D. Centralization Exposure with Control Variables
Cohort’s exposure to centralization 2.03 0.64 1.73 2.51 1.73 0.60 0.43
  $\times$ Tokyo area ($<$100 km) (0.001)*** (0.008)*** (0.020)** (0.002)*** (0.024)** (0.006)*** (0.074)*
[0.005]*** [0.001]*** [0.007]*** [0.004]*** [0.055]* [0.017]** [0.188]
Observations 705 705 705 705 705 705 705
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 8.77 1.24 5.19 6.21 7.99 1.34 0.82
Mean dep var (Tokyo area 10.62 1.51 6.75 6.86 11.21 1.25 1.02
   under decentralization)
[…]
Table 3: Long-run Impacts of Centralization: Destinations of Career Elites
(1) (2) (3) (4) (5) (6) (7)
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors
A. Baseline Specification
Age 17 under centralization 4.20 0.48 2.03 3.11 3.47 0.61 0.71
  $\times$ Tokyo area ($<$100 km) (0.060)* (0.151) (0.320) (0.056)* (0.238) (0.000)*** (0.120)
[0.000]*** [0.154] [0.060]* [0.005]*** [0.003]*** [0.077]* [0.092]*
B. Adding Control Variables
Age 17 under centralization 3.16 0.74 2.44 2.90 3.64 0.41 0.54
  $\times$ Tokyo area ($<$100 km) (0.021)** (0.221) (0.263) (0.038)** (0.223) (0.017)** (0.043)**
[0.008]*** [0.010]*** [0.012]** [0.010]*** [0.002]*** [0.122] [0.134]
Observations 705 705 705 705 705 705 705
Birth cohort FE, Pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 8.30 1.21 5.08 5.88 7.74 1.27 0.78
Mean dep var (Tokyo area 14.73 2.46 9.60 9.64 15.54 2.14 1.32
   under decentralization)
Notes: This table shows difference-in-differences estimates of the long-run effects of the centralized admission system on the geographical destinations of career elites. By birth cohort (born in 1880–1894), we count the number of elites who reside in each prefecture in 1939. Unlike the previous tables, all outcome variables are measured at the prefecture of residence. In Panel B, we control for time- and cohort-varying prefecture characteristics, i.e., the number of primary schools in the prefecture in the year when the cohort turned age 17, the number of middle-school graduates in the prefecture in the year when the cohort turned age 17, log GDP of the prefecture when the cohort turned age 20, and birth population of the cohort in the prefecture. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain wild cluster bootstrap p-values based on standard errors clustered at the cohort level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.1 for discussions about this table.
Table 4: Long-run Impacts of Centralization: National Production of Top Government Officials
(1) (2) (3) (4)
Exam passers Exam passers Exam passers Exam passers
graduated graduated not
from from graduated
School 1 Schools 2–8 from
(Tokyo) Schools 1–8
Centralized 5.96 -2.81 14.93 -12.11
(0.802) (0.221) (0.014)** (0.044)**
Observations 33 33 33 33
Control time trend Yes Yes Yes Yes
Control exam passers No Yes Yes Yes
Mean dep var 189.55 37.29 68.47 83.80
Mean dep var (decentralization) 194.22 37.96 64.09 92.17
(1) (2) (3) (4)
Top-ranking Top-ranking Top-ranking Top-ranking
officials officials officials officials
graduated graduated not
from from graduated
School 1 Schools 2–8 from
(Tokyo) Schools 1–8
Centralized 4.19 -0.48 5.22 -0.55
(0.006)*** (0.648) (0.000)*** (0.596)
Observations 33 33 33 33
Control time trend & exam passers Yes Yes Yes Yes
Mean dep var 29.77 8.27 12.97 8.53
Mean dep var (decentralization) 28.66 8.30 11.22 9.14
Notes: Panel (a) shows OLS estimates of the effects of the centralized admissions on the number of individuals who passed the administrative division of the Higher Civil Service Exams (administrative HCSE). Panel (b) shows OLS estimates of the effects of the centralized admissions on the number of top-ranking higher civil officials. The estimates are based on the cohort level data (1898–1930), where cohort is defined by the year of entering a higher school or its equivalent. The data is compiled from the complete list of individuals who passed the administrative HCSE in 1894–1941 and their biographical information. “Exam passers” is the number of individuals in cohort $t$ who passed the administrative HCSE. “Top-ranking officials” is the number of top-ranking officials in cohort $t$ (i.e., the number of individuals who entered a higher school or its equivalent in year $t$, passed the administrative HCSE, and were internally promoted to the top three ranks of higher civil service in their lifetime). “Centralized” is the indicator variable that takes 1 if cohort $t$ entered a higher school or its equivalent under the centralized admissions in 1902–07, 1917–18, and 1926–27. “Mean dep var (decentralization)” is the mean of the dependent variable for the cohorts who entered a higher school or its equivalent under the decentralized admissions. In all regressions, we control for quadratic time trends. Parentheses contain P values based on Newey-West standard errors with the maximum lag order of 3. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.2 for discussions about this table.

### v2

@@ 4.2 National Production of Elites
The above analysis examines the distributional consequence of the centralized admissions. A complementary problem is its overall impact for the whole country. To study it, we now turn to whether meritocratic centralization increased the national production of occupational elites in the long run. We first focus on a specific subgroup of elites—higher civil officials—for whom we have complete-count data from administrative records. Top-ranking civil officials are also a key group of occupational elites; in our empirical context, higher civil service was considered to be one of the most prestigious jobs due to their political and legislative influence. As a result, high-achieving students at the top universities competed to enter the Ministry of Finance and other selective ministries (Shimizu, 2019). We investigate whether cohorts exposed to the centralized admissions produced a greater number of top-ranking civil officials compared to cohorts exposed to the decentralized admissions. After that, we also analyze the impact of centralization on the national production of all occupational elites listed in the PIR.
@@
Our main data source is the list of individuals who passed the Higher Civil Service Examinations (HCSE) and their biographical information (Hata, 1981). The HCSE were selective national qualification exams held annually from 1894 to 1947. The 1893 ordinance required all individuals to pass the HCSE for appointment in the administrative division of higher civil service with some exceptions for special appointments (Spaulding, 1967, Chapter 25; Shimizu, 2019, Chapter 5). We digitized the information of all individuals who passed the administrative division of the HCSE in 1894–1941, including their full name, education (both university and higher school), year of university graduation, year of passing the exam, starting position, final position, year of retirement, and other notable positions held.
We define “top-ranking officials” as higher civil officials who were internally promoted to one of the top three ranks by the end of their career. In the prewar Japanese bureaucracy system, the higher civil service refers to the top ten ranks of national government offices in the administrative, judicial, and diplomatic divisions. Within the higher civil service, the top three ranks were distinctively called “imperial appointees.” The first rank consisted of minister level positions, while the second and third ranks consisted of vice minister level positions (such as vice minister, director general, bureau chief, and prefectural governor). Appendix A.4.2 provides more details.
We count the number of top-ranking officials and the number of individuals who passed the administrative division of the HCSE (hereafter “exam passers”) by cohort. Here cohort is defined by “the year of entering a higher school or its equivalent.” Since we only observe the year of university graduation in the data, we estimate “the year of entering a higher school or its equivalent.” We first find the exact year of entering a higher school for all top-ranking officials who graduated from Schools 1–8 by searching their names in Student Registers of Schools 1–8. Using this information, we then estimate “the year of entering a higher school or its equivalent” for the rest of higher civil officials using the method described in Appendix A.4.2.
@@
Effect on Higher Civil Service Exam Passers
As an intermediate outcome, we first examine the impacts of the centralized admission system on the number of individuals who passed the selective HCSE. We divide exam passers into three mutually exclusive subgroups: (a) those who graduated from most selective urban School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates. For each subgroup, we count the number of exam passers by cohort. We estimate the following equation for the entire group and for each subgroup:
@@
where $Y_{t}$ is the number of exam passers in a given group in cohort $t$ (defined by the year of entering a higher school or its equivalent), and $Centralized_{t}$ is the indicator that takes 1 if cohort $t$ entered a higher school or its equivalent during centralization. For a subgroup regression, we control for the total number of exam passers in cohort $t$ (denoted by $X_{t}$). We also control for a quadratic time trend where $Trend_{t}$ is the number of years since 1897.
Centralization changed the composition of exam passers. In Table 4 Panel (a), Column (3) shows that centralization increased the number of exam passers who graduated from Schools 2–8 by 23%, while Column (4) indicates that centralization reduced the number of exam passers who are not Schools 1–8 graduates by 13%. By contrast, centralization did not have a significant effect on the total number of exam passers or the number of exam passers who graduated from School 1 (Columns (1) and (2)). This reflects the fact that the total number of exam passers was determined mainly by demand-side factors (e.g., ministries’ needs) rather than supply-side factors (e.g., the quality of exam takers). Thus, the main impact of centralization was not on the number but on the composition of exam passers, and its positive effect was concentrated on Schools 2–8 graduates. This finding is consistent with the result of our short-run analysis that centralization led to a large increase in Tokyo-area-born high-achieving entrants to Schools 2–8 (Table 1 Panel (b) Columns (2)–(8)), which likely improved the academic standing of graduates of these schools.
@@
We now analyze the impact of centralization on the national production of top-ranking officials. Our main outcome variable is the number of higher civil officials who were internally promoted to the top three ranks by the end of their career. We count the number of top-ranking officials by cohort and run the same regression as above for each group. For all regressions, we control for the total number of exam passers from the cohort.
@@
Meritocratic centralization produced a larger number of high-quality bureaucratic elites for the whole country. Table 4 Panel (b) Column (1) shows that the total number of top-ranking officials increased for centralized cohorts by 15%, with high statistical significance. This result is consistent with the idea that the centralized system (which assigned top-scoring students to Schools 1–8) increased the quality of Schools 1–8 entrants, which in turn improved the likelihood of Schools 1–8 graduates to win promotion competitions over close decentralized cohorts. Indeed, Column (3) indicates that centralization had a large, positive, and significant effect on the number of top-ranking officials who graduated from Schools 2–8, where the number of urban-born high-achieving entrants spiked during centralization. The number of top-ranking officials who graduated from Schools 2–8 increased by 47% for centralized cohorts compared to decentralized cohorts. By contrast, in Column (2), centralization only had a small and insignificant effect on the number of top-ranking officials who graduated from School 1, where the number of urban-born high-scoring entrants did not change much during centralization.
@@
Schools 1–8’s impacts on the likelihood of promotion to top-ranking officials appear especially large for high-achieving students. Indeed, centralization’s positive effect on Schools 2–8 graduates was so large (Column (3)) that it dominated centralization’s small and negative effect on those who did not graduate from Schools 1–8 (Column (4)), resulting in a positive effect on the total number of top-ranking officials (Column (1)). This finding suggests that the meritocratic assignment of high achievers to Schools 1–8 produced a net positive effect on the national production of top-ranking officials. This result is inconsistent with the pure selection hypothesis that the role of the centralized admissions is simply to select and send a fixed number of high-achieving students to national higher education and that the value added of national higher education is homogeneous across all students.
@@
Schools 1–8’s particularly large value-added for high-achieving students can be due to multiple factors. First, human capital returns to education by these schools could be higher for high-achieving students (due to demanding curriculum, for example). Gathering high-achieving students in these schools may also produce higher peer effects. Second, high-achieving students may benefit more from their signaling value or from gaining connections with powerful alumni than low-achieving students. While our data and variation do not allow for a clear decomposition into these different factors, our analysis suggests that the effect of meritocratic centralization remains similar even after controlling for the size of same-cohort colleagues from the same schools or focusing on those with the highest educational qualification (Appendix Section A.4.4). Much of Schools 1–8’s effects therefore likely come from the first human capital factor.
@@
One potential threat to our identification is that the number of available top-ranking positions might increase during the centralization periods. To investigate this concern, recall that we define our cohort by the year of entering a higher school or its equivalent (not by the year of becoming top-ranking officials). Within each cohort, the number of years taken from entering a higher school to the appointment for the first top-ranking position varied widely from 20 to 30 years (see Appendix Figure A.9 and Appendix A.4.3 for details). Individuals in a given cohort were therefore promoted to top-ranking positions in different years. This limits the concern that a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions may drive our results.
@@
Another potential threat is that the number of top-ranking officials may be differentially impacted by wars between decentralized and centralized cohorts. During the US occupation of Japan right after the WWII, a substantial number of top officials were purged from public service, and a few were sentenced to death for war crimes. However, we show that the number of top-ranking officials who died in wars, or were purged or executed after WWII, changes little between decentralized and centralized cohorts (Appendix Table A.17).
@@
The last potential concern is that the number of top bureaucrats might increase at the expense of other types of elites. To examine this possibility, we use the 1939 PIR data to compute the share of top bureaucrats among all socially distinguished individuals listed in the PIR. As Appendix Figure A.10 shows, we observe no positive association between the share of government officials and the cohort’s exposure to the centralized admissions (although the share of government officials is rising over time reflecting a growing public sector). Therefore, our result is unlikely to be driven by students’ brain drain from other elite occupations into top bureaucrats under the centralized admissions.
Effect on All Occupational Elites
Moreover, we find that meritocratic centralization produced more occupational elites across fields, not just government officials. As Appendix Figure A.11 shows, higher exposure to the centralized system is associated with a greater total number of all occupational elites (as defined in Section 4.1) listed in the PIR (1939). A potential concern is that this analysis may conflate age effects and cohort effects because the centralized cohorts, who were in their fifties in the PIR (1939), may have a higher likelihood of holding prominent positions simply due to their age rather than the centralized system’s impact. To distinguish the cohort effect from the age effect, we additionally use the 1934 edition of the PIR, in which we observe the centralized cohorts in their late forties. The total number of all occupational elites listed in the PIR (1934) also increased for the centralized cohorts, as Appendix Figure A.11 shows.
Table 5 shows this result in a regression analysis using both 1934 and 1939 editions of the PIR. We count the number of individuals who are aged 40–69 by birth cohort in each PIR edition, and obtain the cohort-edition level data. We regress this count on each cohort’s exposure to centralization, while controlling for non-linear age trends defined by either quadratic age trends, quartic age trends, or edition-specific quadratic age trends. We also control for the edition fixed effect. Panel A shows that the total number of all occupational elites listed in the PIR is 5–13% higher among centralized cohorts. Looking at some specific groups of elites, the number of top 0.05% income earners and the number of medal recipients increased by 8–10% (Panel B) and 12-24% (Panel C), respectively, among centralized cohorts. The centralized system may therefore be more productively efficient not only for producing top bureaucrats but also for producing other types of individuals in leadership positions.
@@ 5 Conclusion
The design of selective school admissions persistently impacts the production and distribution of society’s leaders. We reveal this fact by looking at the world’s first recorded use of nationally centralized admissions and its subsequent abolitions in the early twentieth-century. While centralization was designed to make the school seat allocations more meritocratic, there turns out to be a tradeoff between meritocracy and equal access to selective higher education and career advancement. Meritocratic centralization led students to apply to more selective schools and make more inter-regional applications. As high-achieving students were located disproportionately in urban areas, however, centralization caused urban high achievers to crowd out rural applicants from advancing to higher education.
@@
Most importantly, these impacts persist: Several decades later, the centralized system produced more individuals in leadership positions, especially top-ranking government officials. The distributional effect also persists: Meritocratic centralization increased the number of high income earners, medal recipients, and other occupational elites coming from urban areas relative to those from rural areas.
@@
Even though our study uses the admission reforms unique to our historical setting, the implications of our study may have broader relevance. For instance, distributional consequences of centralized meritocratic admissions may be a reason why many countries continue to use seemingly inefficient decentralized college admissions. While this paper focuses on the particular decentralized system, the takeaway from our results would be relevant for other types of decentralized admissions as long as costly applications make applicants self-select into a limited number of schools. For any such decentralized systems, the meritocracy-equity tradeoff is an important concern.
@@
Methodologically, the use of natural experiments in history may be also valuable for studying the long-run effects of market designs in other areas, such as housing, labor, and health markets. The disadvantage of using historical events is the limited availability of data. The ideal way to alleviate the data concerns would be to use modern administrative data. One may imagine linking administrative tax return data and school district data to measure the long-run effects of school choice reforms around the world in the past few decades. Such an effort would be a fruitful complement to our historical study.